OpenStudy (anonymous):

What is the initial value of the equation shown? y = −2x − 5 −5 −2 2 5

OpenStudy (anonymous):

x=6 month y=12 month x+y=125 112x + 92y=12660 -92x-92y=-11500, multiplying the top by -92 112x+92=12660 20x=1160 x=58 6 month CD y=67 12 month CD 58*112=$6496 67*92=$6194 sum is $12,660

OpenStudy (anonymous):

okay what

OpenStudy (anonymous):

lol u still dont get it

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

no and I really need help with this test or I will fail the class

OpenStudy (p0sitr0n):

If by initial value you mean y intercept, then take x=0. Then \[y=-2(0)-5\ \implies y=0-5 \implies y=-5\]

OpenStudy (p0sitr0n):